# Noise2Score3D: Tweedie’s Approach for Unsupervised Point Cloud Denoising

### v2

@@ 1 Introduction
Inspired by this, we propose an unsupervised point cloud denoising method, Noise2Score3D, that effectively denoise point clouds using only noisy point clouds as input. Our method achieves state-of-the-art performance among unsupervised methods and offers improved generalization ability across different datasets, as well as noise levels without re-training. Noise2Score3D also achieves high efficiency, as denoising is performed in one step with known noise parameters. The main contributions of this paper are as follows:
@@
  • Combining score estimation with Tweedie’s formula, we propose an unsupervised learning framework for point cloud denoising. Experiments demonstrate that, Noise2Score3D achieves state-of-the-art results on several benchmark datasets with varying noise levels in terms of both quantitative metrics and visual quality.
@@ 2.3 Unsupervised image denoising
Figure 2: Illustration of different classes of learning-based methods. Hollow circles represent the initial positions of data points, while filled circles indicate adjusted positions.
@@ 3 Method
Our proposed framework, e.g. Noise2Score3D, adopts a Bayesian approach to unsupervised point clouds denoising. Specifically, Noise2Score3D consists of two steps: 1) estimating the score function from the noisy point cloud, and 2) restoring the clean point cloud using the estimated score and noise parameters. This decoupling between model estimations and denoising (displacements calculation) allows Noise2Score3D to have unique advantages over other unsupervised learning methods, the most important of which is that the same loss function can be used regardless of the noise model and parameters. By utilizing Tweedie’s formula from Bayesian statistics, the denoising process can be done in one step, given the noise parameters. These features distinguish our methods from other unsupervised denoising methods, such as TotalDn [21] and ScoreDenoise [36]. Fig. 2 illustrates the differences between our method and previous displacement or sampling-based learning approaches. As a result, our approach is simple to train and use, and does not expect the user to provide parameters to characterize the surface or noise model. In Sec. 3.1, we first introduce the basic theory. In Sec. 3.2, we elaborate on each proposed step, including the score estimation framework with the corresponding network structure and loss function. Finally, in Sec. 3.3, we describe the denoising process, emphasizing on the case of unknown noise parameters.
@@ 3.2 Score estimation
Network for score estimation We build upon the Kernel Point Convolution (KPConv) architecture introduced by Thomas et al. [47] for score estimation. KPConv network is a point cloud feature encoder which accurately captures local geometric structures and adapts to varying point densities. We modified it to accept only the three-dimensional coordinates of the noisy point cloud $y$ as input features. The convolution weights of KPConv are determined by Euclidean distances to kernel points, and the number of kernel points is set to 15 in our implementation. Note that the radius neighborhood (e.g., increasing from 0.1 to 1.6 across stages) is used to maintain a consistent receptive field, while grid subsampling is applied in each layer to achieve high robustness under varying densities of point clouds. Our network comprises five encoder stages with a total of 14 blocks: an initial KPConv block and a residual block in the first stage, followed by four stages, each with three residual blocks outputting 256, 512, 1024, and 2048 channels, respectively. The decoder includes four blocks, reducing the feature dimension from 3072 to 128, and a final fully connected layer outputs three-dimensional vectors (estimated score values) for each point. The network architecture is illustrated in Fig. 1 with total parameter-number of 24.3M. The output score values are interpreted as the values statistically closest to the gradient of the log-probability density function $S(y)$ at each point. The estimated score function is used for reconstructing the clean point cloud during the denoising stage, as detailed in the later part of Sec. 3.1.
@@ 3.3 Denoising with unknown noise parameters
where $\hat{x}(\sigma)=y+\sigma^{2}S(y)$. This allows us to automatically find the best noise parameter from the data, ensuring that the processed point cloud achieves optimal smoothness while minimizing geometric distortion. The pseudo code for denoising using $TV_{PC}$ is shown in Algorithm 1.
@@ 4 Experiments
Figure 3: Visualization results by different algorithms on synthetic datasets: (a) ModelNet-40 with 50k points; (b) PU-Net with 10k points; (c) PU-Net with 50k points; (d) ModelNet-40 with simulated LiDAR noise. All shown results are with noise level $\sigma=2\%$. Points with smaller error are colored more blue, and otherwise colored yellow.
@@ 4.2 Quantitative Results
We test our trained model on three types of datasets, including the ModelNet-40 dataset/ PU-Net dataset with isotropic Gaussian noise [55], and ModelNet-40 with simulated LiDAR noise [21]. For Gaussian noises, the noise level (standard deviation of Gaussian) is set to 1%, 2%, and 3% of the radius of the shape boundary sphere. Below we report the results for each case. For all results, we assume the noise parameter is unknown and is estimated following the procedure in Sec. 3.3.
Accuracy resultsTab. 1 shows the comparative performance of various unsupervised denoising methods across sparse (10k) and dense (50k) point clouds with the ModelNet-40 dataset. It can be observed that Noise2Score3D consistently achieves better results than other unsupervised approaches in both Chamfer distance (CD) and point-to-mesh (P2M) metrics, and rivals GLR, the best-performing classical method. Notably, our method demonstrates consistent performance across various noise levels, particularly at high noise levels where other unsupervised learning methods may produce large errors.
To assess the model’s generalization ability, we reevaluate it on the PU-Net dataset (Tab. 2). It can be seen that Noise2Score3D outperforms other unsupervised learning-based methods, highlighting its strong generalization ability across datasets. It is worth noting that our model is trained solely on the ModelNet-40 dataset, whereas the classical methods and learning-based methods are optimized/trained on the PU-Net dataset directly. We also tested the accuracy of the inference results using the true noise parameters and found that CD and P2M could be further improved (for details please refer to the Supplementary Material Tab. 7).
We also conducted evaluations on a synthetic dataset with emulated LiDAR sensor noise [21]. The results (Table 3) show that Noise2Score3D significantly outperforms other unsupervised learning methods across various noise levels and approaches the performance of the best classical method, GLR. Overall, our method demonstrates persistent denoising performance across various noise conditions.
Inference time We compare the inference time of Noise2Score3D with other learning-based methods (TotalDn, DMR-U and Score-U). Experiments are run on the same GPU device (NVIDIA RTX 3080ti) with the exception for TotalDn, which we can only run on a NVIDIA RTX 2080Ti GPU. The classical methods are not compared here for their different implementation platforms (Matlab). We report the average inference time on 20 PU-Net point clouds with 10K or 50K sampled points. As shown in Tab. 4, Noise2Score3D is, on average, 15 (50) or more times faster in inference times than other methods on 10K (50K) point clouds, confirming the efficiency of our approach.
@@ 4.3 Qualitative Results
We visualize denoising results of the proposed method and the competing baselines under Gaussian and simulated LiDAR noise, as shown in Fig. 3. The color of each point indicates its reconstruction error, which is measured only by the point-to-mesh distance, unlike in Sec. 4.1, where P2M consists of two distances: point-to-mesh and mesh-to-point. From Fig. 3, it is evident that our results are cleaner and visually superior compared to those of other unsupervised learning methods and most classical approaches. Furthermore, we conduct a qualitative study on the real-world dataset Paris-rue-Madame[44]. The results are shown in Fig. 4, in which the left and right parts of each sub-figure are rendered from different views. Due to the varying and intricate nature of noise in point clouds, many methods struggle to maintain a balance between effective denoising and the preservation of fine details, particularly under high-conditions. Additionally, point clouds produced by these methods often suffer from nonuniform distribution, as shown in the right part of the subfigures in Fig. 4. In contrast, Noise2Score3D produces cleaner and smoother results, in the meanwhile better preserving the details.
@@ 4.4 Ablation Studies
We performed two ablation studies:
Choices of loss function We compare the current results of AR-DAE loss with that of using DAE loss [1] in Tab. 5, showing the superiority of AR-DAE loss. Training and evaluation are both on the ModelNet-40 dataset with 50k points.
@@
Annealing noise levels $\sigma_{t}$ in training We compare the results of the current model trained by gradually annealing noise levels $\sigma_{t}$ with those of a fixed model $\sigma_{t}$. The results are shown in Tab. 6, from which we can see that, while the model trained with fixed $\sigma_{t}$ can get better results for test data with similar noise levels, the annealing procedure enables stable denoising outcomes across different noise levels.
@@
Table 6: Effect of different $\sigma_{t}$ choices in training.
@@ 5 Conclusions
The current method is restricted to exponential family noises, and assumes an uniform noise level across all points in the cloud. But it is possible to be extended to more complex noise models, such as multiplicative and structurally correlated noise[52]. Future work includes extending the method to other scenarios, for example, color point clouds and more realistic noise setup. Another direction is to apply the score estimation framework to other point cloud related tasks, e.g. point cloud upsampling and completion. We believe that our approach can inspire future research in the application of Bayesian methods in point cloud denoising and, more generally, 3D data processing.
@@ A.1 Definition of Total Variation for Point Cloud
A.2 Validation of $TV_{PC}$
Figure 5: Change of average $TV_{PC}$, $CD$ and $P2M$ values with noise parameter $\sigma$ using the PU-Net dataset[55] .
PU-Net 10K 50K
Inference 1% 2% 3% 1% 2% 3%
CD P2M CD P2M CD P2M CD P2M CD P2M CD P2M
Table 7: Comparison of denoising results using estimated noise parameters by $TV_{PC}$ ($\sigma_{tv}$) vs. true values ($\sigma_{true}$) on the PU-Net dataset[55] .
Fig. 5 shows the $TV_{PC}$, $CD$ and $P2M$ values with different noise parameter $\sigma$ for the inference result on PU-Net dataset with $\sigma$=0.03. We can observe that $TV_{PC}$ follows a similar trend as $CD$ and $P2M$ and reaches a minimal around the true $\sigma$ value. Note that the clean point clouds are not needed for calculation of $TV_{PC}$. This validates the use of $TV_{PC}$ in estimating unknown noise parameters for real world datasets.
Additionally, Tab. 7 shows the results with noise parameter determined by the proposed $TV_{PC}$ versus that using the true value of $\sigma$. We observe that the metrics of denoising outcomes are similar between two cases across different noise levels, validating the proposed $TV_{PC}$ in assessing the quality of denoising result and estimating the unknown noise parameters.
B Additional visualization results
Figs. 6, 7, 8, 12, 13, 14, 9, 10 and 11 provide additional visualization results on ModelNet-40 and PU-Net. Points with smaller error are colored more blue, and otherwise colored yellow. Note that Fig. 3 in the main article shows results with the noise scale set to 2% of the bounding sphere’s radius for the Gaussian noise and 1.5% for the simulated LiDAR noise.

### v3

@@ 1 Introduction
Inspired by this, we propose Noise2Score3D, an unsupervised point cloud denoising method that, using only noisy point clouds as input, predicts the score function and applies the Tweedie’s formula to estimate the clean point positions based on the predicted score. Our method achieves state-of-the-art performance among unsupervised methods and offers improved generalization ability across different datasets, as well as noise levels without re-training. Noise2Score3D also achieves high efficiency, as denoising is performed in one step with known noise parameters. Our source-code is publicly available at here.
The main contributions of this paper are as follows:
@@
  • Combining score estimation with Tweedie’s formula, we propose an unsupervised learning framework for point cloud denoising. Experiments demonstrate that Noise2Score3D achieves state-of-the-art results on several benchmark datasets with varying noise levels in terms of both quantitative metrics and visual quality.
@@ 3 Method
Our proposed framework, e.g. Noise2Score3D, adopts a Bayesian approach to unsupervised point clouds denoising. Specifically, Noise2Score3D consists of two steps: 1) estimating the score function from the noisy point cloud, and 2) restoring the clean point cloud using the estimated score and noise parameters. This decoupling between model estimations and denoising (displacements calculation) allows Noise2Score3D to have unique advantages over other unsupervised learning methods, the most important of which is that the same loss function or pre-training weight can be used regardless of the noise model and parameters. By utilizing Tweedie’s formula from Bayesian statistics, the denoising process can be done in one step, given the noise parameters. These features distinguish our methods from other unsupervised denoising methods, such as TotalDn [21] and ScoreDenoise [36]. As a result, our approach is simple to train and use, and does not expect the user to provide parameters to characterize the surface or noise model. In Sec.˜3.1, we first introduce the basic theory. In Sec.˜3.2, we elaborate on each proposed step, including the score estimation framework with the corresponding network structure and loss function. Finally, in Sec.˜3.3, we describe the denoising process, emphasizing on the case of unknown noise parameters.
@@ 3.2 Score estimation
Network for score estimation We build upon the Kernel Point Convolution (KPConv) architecture introduced by Thomas et al. [47] for score estimation. KPConv network is a point cloud feature encoder which accurately captures local geometric structures and adapts to varying point densities. We modified it to accept only the three-dimensional coordinates of the noisy point cloud $y$ as input features. The convolution weights of KPConv are determined by Euclidean distances to kernel points, and the number of kernel points is set to 15 in our implementation. Note that the radius neighborhood (e.g., increasing from 0.1 to 1.6 across stages) is used to maintain a consistent receptive field, while grid subsampling is applied in each layer to achieve high robustness under varying densities of point clouds. Our network comprises five encoder stages with a total of 14 blocks: an initial KPConv block and a residual block in the first stage, followed by four stages, each with three residual blocks outputting 256, 512, 1024, and 2048 channels, respectively. The decoder includes four blocks, reducing the feature dimension from 3072 to 128, and a final fully connected layer outputs three-dimensional vectors (estimated score values) for each point. The network architecture is illustrated in Fig.˜1 with total parameter-number of 24.3M. The output score values are interpreted as the values statistically closest to the gradient of the log-probability density function $S(y)$ at each point. The estimated score function is used for reconstructing the clean point cloud during the denoising stage, as detailed in the later part of Sec.˜3.1.
@@ 3.3 Denoising with unknown noise parameters
where $\hat{x}(\sigma)=y+\sigma^{2}S(y)$. This allows us to automatically find the best noise parameter from the data, ensuring that the processed point cloud achieves optimal smoothness while minimizing geometric distortion. The pseudo code for denoising using $TV_{PC}$ is shown in Algorithm˜1.
@@ 4 Experiments
Figure 2: Visualization results by different algorithms on synthetic datasets: (a) ModelNet-40 with 50k points; (b) PU-Net with 10k points; (c) PU-Net with 50k points; (d) ModelNet-40 with simulated LiDAR noise. Note that shown results with the noise scale set to 2% of the bounding sphere’s radius for the Gaussian noise and 1.5% for the simulated LiDAR noise. Points with smaller error are colored more blue, and otherwise colored yellow.
@@ 4.2 Quantitative Results
We test our trained model on three types of datasets, including the ModelNet-40 dataset/ PU-Net dataset with isotropic Gaussian noise [55], and ModelNet-40 with simulated LiDAR noise [21]. For Gaussian noises, the noise level (standard deviation of Gaussian) is set to 1%, 2%, and 3% of the radius of the shape boundary sphere. Below we report the results for each case. For all results, we assume the noise parameter is unknown and is estimated following the procedure in Sec.˜3.3.
Accuracy resultsTab.˜1 shows the comparative performance of various unsupervised denoising methods across sparse (10k) and dense (50k) point clouds with the ModelNet-40 dataset. It can be observed that Noise2Score3D consistently achieves better results than other unsupervised approaches in both Chamfer distance (CD) and point-to-mesh (P2M) metrics, and rivals GLR, the best-performing classical method. Notably, our method demonstrates consistent performance across various noise levels, particularly at high noise levels where other unsupervised learning methods may produce large errors.
To assess the model’s generalization ability, we reevaluate it on the PU-Net dataset (Tab.˜2). It can be seen that Noise2Score3D outperforms other unsupervised learning-based methods, highlighting its strong generalization ability across datasets. It is worth noting that our model is trained solely on the ModelNet-40 dataset, whereas the classical methods and learning-based methods are optimized/trained on the PU-Net dataset directly. We also tested the accuracy of the inference results using the true noise parameters and found that CD and P2M could be further improved (for details please refer to the Supplementary Material Tab.˜7).
We also conducted evaluations on a synthetic dataset with emulated LiDAR sensor noise [21]. The results (Table 3) show that Noise2Score3D significantly outperforms other unsupervised learning methods across various noise levels and approaches the performance of the best classical methods. Overall, our method demonstrates persistent denoising performance across various noise conditions.
@@ 4.3 Qualitative Results
We visualize denoising results of the proposed method and the competing baselines under Gaussian and simulated LiDAR noise, as shown in Fig.˜2. The color of each point indicates its reconstruction error, which is measured only by the point-to-mesh distance, unlike in Sec.˜4.1, where P2M consists of two distances: point-to-mesh and mesh-to-point. From Fig.˜2, it is evident that our results are cleaner and visually superior compared to those of other unsupervised learning methods and most classical approaches. Furthermore, we conduct a qualitative study on the real-world dataset Paris-rue-Madame[44]. The results are shown in Fig.˜3, in which the left and right parts of each sub-figure are rendered from different views. Due to the varying and intricate nature of noise in point clouds, many methods struggle to maintain a balance between effective denoising and the preservation of fine details, particularly under high-conditions. Additionally, point clouds produced by these methods often suffer from nonuniform distribution, as shown in the right part of the subfigures in Fig.˜3. In contrast, Noise2Score3D produces cleaner and smoother results, in the meanwhile better preserving the details.
4.4 Inference time
We compare the inference time of Noise2Score3D with other learning-based methods (TotalDn, DMR-U and Score-U). Experiments are run on the same GPU device (NVIDIA RTX 3080ti) with the exception for TotalDn, which we can only run on a NVIDIA RTX 2080Ti GPU. The classical methods are not compared here for their different implementation platforms (Matlab). We report the average inference time on 20 PU-Net point clouds with 10K or 50K sampled points. As shown in Tab.˜4, Noise2Score3D is, on average, 15 (50) or more times faster in inference times than other methods on 10K (50K) point clouds, confirming the efficiency of our approach.
@@ 4.5 Ablation Studies
Choices of loss function We compare the current results of AR-DAE loss with that of using DAE loss [1] in Tab.˜5, showing the superiority of AR-DAE loss. Training and evaluation are both on the ModelNet-40 dataset with 50k points.
@@
Annealing noise levels $\sigma_{t}$ in training We compare the results of the current model trained by gradually annealing noise levels $\sigma_{t}$ with those of a fixed model $\sigma_{t}$. The results are shown in Tab.˜6, from which we can see that, while the model trained with fixed $\sigma_{t}$ can get better results for test data with similar noise levels, the annealing procedure enables stable denoising outcomes across different noise levels.
Table 6: Effect of different $\sigma_{t}$ choices in training.
@@
Result comparison using noise parameters estimated with $TV_{PC}$ vs. true valuesTab.˜7 shows the results with noise parameter determined by the proposed $TV_{PC}$ versus that using the true value of $\sigma$. We observe that the metrics of denoising outcomes are similar between two cases across different noise levels, validating the proposed $TV_{PC}$ in assessing the quality of denoising result and estimating the unknown noise parameters.
Table 7: Denoising results using estimated noise parameters by $TV_{PC}$ ($\sigma_{tv}$) vs. true values ($\sigma_{true}$) on the PU-Net dataset[55].
PU-Net 10K 50K
Inference 1% 2% 3% 1% 2% 3%
CD P2M CD P2M CD P2M CD P2M CD P2M CD P2M
@@ 5 Conclusions
The current method is restricted to exponential family noises, and assumes an uniform noise level across all points in the cloud. But it is possible to be extended to more complex noise models, such as multiplicative and structurally correlated noise[52]. Future work includes extending the method to other input types, e.g. color point clouds and more realistic noise setup. Another direction is to apply the score estimation framework to other tasks, e.g. point cloud upsampling and completion. We believe that our approach can inspire future research in the application of Bayesian methods in point cloud denoising and, more generally, 3D data processing.
Acknowledgement
The work was supported by Guangdong Provincial Quantum Science Strategic Initiative (GDZX2306001, GDZX2303001) the startup fund of Shenzhen City, Shenzhen Fundamental research project (Grant No. JCYJ20241202123903005).
@@ A.1 Definition of Total Variation for Point Cloud
A.2 Validation of $TV_{PC}$ in estimating the noise level
Fig.˜4 shows the $TV_{PC}$, $CD$ and $P2M$ values with different noise parameter $\sigma$ for the inference result on PU-Net dataset with $\sigma$=0.03. We can observe that $TV_{PC}$ follows a similar trend as $CD$ and $P2M$ and reaches a minimal around the true $\sigma$ value. Note that the clean point clouds are not needed for calculation of $TV_{PC}$. This validates the use of $TV_{PC}$ in estimating unknown noise parameters for real world datasets.
Figure 4: Change of average $TV_{PC}$, $CD$ and $P2M$ values with noise parameter $\sigma$ on the PU-Net dataset[55] .
B Performance Comparison and Ablation Study of Point Cloud Backbones
Tab.˜8 evaluates the effectiveness of different point cloud backbones,the analysis reveals that the KPConv backbone does not inherently improve the performance of DMR or Score-U, suggesting that the superior results of our method are not solely attributable to the KPConv network architecture but rather to our novel Bayesian approach. Additionally, we show the result using PointNet++ backbone, which underperforms in sparse point cloud regions and under higher noise conditions compared to KPConv, reinforcing the appropriateness of our chosen backbone for the proposed framework.
Table 8: Results on the PU-Net 10k dataset (results of 50k dataset are not included due to space limitation). Results with * are taken from Tab.˜2. in the main text for comparison purpose.
Noise level 1% 2% 3%
Dataset Model CD P2M CD P2M CD P2M
PU-Net DMR-Ukpconv 33.89 28.16 25.53 20.68 20.69 15.99
*DMR-U 5.313 2.522 6.455 3.317 8.134 4.647
Score-ukpconv 3.494 1.077 4.928 1.909 7.250 3.596
*Score-U 3.107 0.888 4.675 1.829 7.225 3.762
OursPointNet++ 3.689 1.287 7.938 4.370 13.382 9.028
*Ours 2.848 1.106 4.190 1.818 5.583 2.947
C Additional visualization results
Figs.˜5, 6, 7, 11, 12, 13, 8, 9 and 10 provide additional visualization results with varying noise levels on ModelNet-40 and PU-Net. Points with smaller error are colored more blue, and otherwise colored yellow.